OpenStudy (anonymous):

(A) You are given a bottle that contains 4.59 cm3 of a metallic solid. The total mass of the bottle and solid is 35.66 g. The empty bottle weighs 14.23 g. What is the density of the solid? (B) mercury is traded by the "flask" a unit has a mass of 34.5 kg. What is the volume of a flask of mercury if the density of mercury is 13.5 g/ML? (C) a thief plans to steal a gold sphere with a radius of 28.9 cm from a museum. If the gold has a density of 19.3 g/cm3 what is the mass of the sphere? [the volume of sphere is V= (4/3)(3.14r3]

OpenStudy (anonymous):

A) 4.67x10^3 kg m^-3

OpenStudy (anonymous):

b) The initial mass is 34.5 kg = 3450g

OpenStudy (anonymous):

How did you get that answer? Please explain it to me. :3

OpenStudy (anonymous):

c) 600mL * (1.28g / mL) = 768g of battery acid

OpenStudy (anonymous):

i am using a graphing calculator so the values might hve +-2 difference

OpenStudy (anonymous):

Ohh. I understand. Thank you so much :) it's really great help.

OpenStudy (anonymous):

that A

OpenStudy (anonymous):

sorry about horrible writing

OpenStudy (anonymous):

for b it is The initial mass is 34.5 kg = 3450g Volume = 3450g * (mL / 13.6g) = 253.68mL

OpenStudy (anonymous):

and C is 600mL * (1.28g / mL) = 768g of battery acid 768g battery acid * (38.1g sulfuric acid / 100 battery acid) = 292.6g sulfuric acid

OpenStudy (anonymous):

Haha. :) that's fine. Actually, we have the same writings. ^_^

OpenStudy (anonymous):

ignore C i gave the wrong

OpenStudy (anonymous):

By the way, what's the solution in letter B?

OpenStudy (anonymous):

for b it is The initial mass is 34.5 kg = 3450g Volume = 3450g * (mL / 13.6g) = 253.68mL <------this

OpenStudy (anonymous):

Thank you qazd. ^_^

OpenStudy (anonymous):

thats the thief one dont ask me from there i got it :)

OpenStudy (anonymous):

Ahh. :) thank you so much qazd. hehe